Over the last 3 evenings, Linda received a total of 64 phone calls at the call center. The second evening, she received 8 more calls than the first evening. The third evening, she received 2 times as many calls as the second evening. How many phone calls did she receive each evening?
Found 2 solutions by mananth, josgarithmetic:
Answer by mananth(16946)

total calls received on 3 days =64
First day we assume she got x calls
On second day she received x+8 calls
Third she received 2(x+8) calls
Therefore x+ x+8 +2(x+8) =64
x+x+8 +2x+16 = 64
4x +24 =64
4x = 40
x=10
Day 1 10 calls
Day 2 10+8 =18 calls
Day 3 18*2 = 36 calls

Answer by josgarithmetic(39623)

EVENING NUMBER        CALLS COUNT
        1                  x
        2                  x+8
        3                 2(x+8)
TOTAL                       64

As uncomplicated as can be.